OpenStudy (anonymous):

A pretzel burned in a calorimeter transfers 15,200 joules to 200.0 g of water. What is the rise in the water’s temperature?

OpenStudy (michele_laino):

here we have to apply the conservation of energy. So, we can write this: \[\Large E = c\;m\;\Delta t\] solving for \(\Delta t\), we get: \[\Large \Delta t = \frac{E}{{c\;m}} = \frac{{15.2}}{{4.18 \cdot 0.2}} = ...^\circ C\] where \(m=200\)grams=\(0.2\)Kg, and \(c\) is the specific heat for water: \(c=4.18 \)kJ/(Kg °C), finally \(E=15,200\) joules= \(15.2\) kJoules